The amounts of money requested on home loan applications at Down River Federal Savings follow the normal distribution, where the amount requested for home loans followed the normal distribution with a mean of $62,000 and a standard deviation of $20,000. (Round z-score computation to 2 decimal places and the final answers to the nearest whole dollars.)

a. What is the minimum amount requested on the largest 7% of loans?

$

b. What is the maximum amount requested on the smallest 13% of loans?

$
Answer #1

Part a)


P ( X > ? ) = 1 - P ( X < ? ) = 1 - 0.07 = 0.93
Looking for the probability 0.93 in standard normal table to calculate critical value Z = 1.48

1.48 = ( X - 62000 ) / 20000
X = 91600
P ( X > 91600 ) = 0.07

Part b)


P ( X < ? ) = 13% = 0.13
Looking for the probability 0.13 in standard normal table to calculate critical value Z = -1.13

-1.13 = ( X - 62000 ) / 20000
X = 39400
P ( X < 39400 ) = 0.13